# TropNNC: Structured Neural Network Compression Using Tropical Geometry

### v2

@@ 5.2.2 Iterative Algorithm for multi-output network
7:  for iter=1 to $\text{iter}=1\text{ to }num\_iter$ do
8:     for $k\in[K]$ do
9:        for $j\in[m]$ do
10:           $\widetilde{C}_{jk}\leftarrow\frac{\langle\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i}),(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\rangle}{\|(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\|^{2}}$
11:        end for
12:        $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
13:     end for
14:  end for
15:  Construct the new linear layer:
16:   For the input weights, the k-th row of weight-bias matrix is $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})$
17:   For the output weights, the k-th column is $\widetilde{\mathbf{C}}_{:,k}^{T}$
Proposition 14
Suppose the clusters K are enough so that for every cluster, no two $(\mathbf{a}_{i}^{T},b_{i}),(\mathbf{a}_{i^{\prime}}^{T},b_{i^{\prime}})$ of the cluster form an obtuse angle. Then, a variant (see Appendix) of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|v(\mathbf{x})-\tilde{v}(\mathbf{x})\|_{1}\leq\sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
6 Extensions, Limitations, and Applications of the Method
Before proceeding to our experiments, we would like to discuss the applicability of our algorithm, highlight some limitations, and explore potential strategies to overcome these challenges.
Deep and Convolutional Networks. Our methods were analyzed previously for the case of networks with a single hidden linear layer with ReLU activations. However, they can be generalized both for deep and convolutional neural networks. In the first case, where we have a deep network with multiple layers, we can view each layer as an instance for our algorithm, by having an input from the previous layer and providing an output to the next layer. Therefore, we can recursively apply our multi-output algorithm to compress each hidden layer of the network.
In the case of convolutional networks, our methods can be extended for convolutional layers with ReLU and maxout or average pooling layers in between them. Here, instead of neurons, we have channels, and instead of weights that connect the neurons, we have kernels that connect the channels. For each hidden layer, the input and output weight matrices $\mathbf{A}$ and $\mathbf{C}$ are 4-dimensional, where each element of the row $(\mathbf{a}_{i}^{T},b_{i})$ and the column $\mathbf{C}_{:,i}$ is a kernel. Compressing a convolutional layer in a structured manner involves clustering channels. To apply our multi-output algorithm, we first flatten the matrices $\mathbf{A}$ and $\mathbf{C}$ so that they become 2-dimensional, by unravelling the kernels row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$. We then apply our multi-output algorithm on the flattened matrices. Finally, after obtaining the new $\widetilde{\mathbf{A}}$ and $\widetilde{\mathbf{C}}$, we reshape the matrices to restore their original dimensions as convolutional layer weight matrices. We emphasize that we are the first to apply a tropical geometrical method to convolutional layers.
Batch normalization and other limitations. Our algorithm can not be applied in the case of networks with batch normalization layers in between the convolutional and linear layers. The reason for this is that batch normalization layers have a non-negligible effect on the zonotopes of the network we approximate. Here, we explain two methodologies that can be used to adapt networks with batch normalization to be applicable for our algorithm. These methodologies were used in the experiments to convert the VGG and ResNet models to equivalent models that our algorithm can be applied to. We propose two methodologies that allow most networks with batch normalization layers to be expressed equivalently as networks without batch normalization. As a first idea, we can fuse together a pair (convolutional (or linear) layer, batch normalization layer) into a single convolutional layer as follows: We take the pair we wish to fuse and form a network with just these two layers. We form a second network with a single, new convolutional layer. We then train the second network to give the same output as the first network on random noise input. This should produce a new convolutional layer that approximates the pair (convolutional (or linear) layer, batch normalization layer) with very little accuracy loss. Finally, we replace each such pair with its approximate convolutional (or linear) layer, which gives us an approximately equivalent network without batch normalization. An alternative second method is to shift the computations of the normalization layer directly to the convolutional or linear layer. Specifically, we can view the batch normalization layer being applied as a multiplication with corrective terms and addition with an extra bias term. We can incorporate these computations into the weights and bias of the convolutional or linear layer.
Using these ideas, we can fuse a batch normalization layer with its preceding convolutional or linear layer. This enables two approaches for applying our algorithm to networks with batch normalization. The first approach involves fusing all layers of the network beforehand, resulting in a fully fused network without batch normalization. The second approach performs fusion during the compression process, applying it layer-by-layer as each layer is compressed. For the second approach, we made a notable observation during our experiments, although the reasoning behind it remains unclear: using the pre-fusion weights/kernels of the layers for clustering vectors consistently led to significantly better results.
As for other limitations, we should note that our algorithm is designed to work strictly on layers as described above. Namely, convolutional or linear layers with ReLU and maxout or average pooling layers in between. Layers that apply computations, like batch normalization layers or the loops of residual blocks of ResNet networks fundamentally change the theory for the analysis of the network. For example, similarly to ThiNet, our methods can not be applied for the compression of the terminal convolutional layer of residual blocks.
@@ 6 Extensions, Limitations, and Applications of the Method
Non-uniform compression. The techniques we have presented apply a uniform pruning ratio to all layers of the network. However, they can easily be extended to non-uniform pruning if instead of K-means we use hierarchical clustering with a global threshold parameter, like in the CUP framework (Duggal et al., 2021). In fact, when used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry. We also propose a modification to step 2 of CUP, specifically the hierarchical clustering step using a global distance threshold. Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
@@
As demonstrated in our experiments, the first variant excels when minimizing the number of parameters, while the second variant is more effective when our objective is to minimize floating-point operation demands.
@@ 7 Experiments
We conduct experiments targeting the compression of the linear and convolutional layers of both convolutional and deep neural networks. Our method TropNNC performs layer-by-layer compression of both linear and convolutional layers and is suitable for regression and classification tasks. The empirical results from these experiments demonstrate that our algorithm achieves the expected performance as claimed. Throughout our experiments, we reserve the use of the iterative variant of TropNNC for cases where its slower performance is justified by the expectation of significantly better results compared to the non-iterative variant.
Baselines. We compare our framework with state-of-the-art methods that conduct structured pruning without requiring re-training (or fine-tuning). Specifically, we compare it with the algorithms proposed by Misiakos et al. (2022), ThiNet (Luo et al., 2017), CUP (Duggal et al., 2021) and the simple baselines random and L1 structured pruning. The algorithms presented in (Misiakos et al., 2022) are originally designed for linear layers only. To enable a fair comparison in compressing convolutional layers, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. ThiNet employs a greedy criterion to remove neurons and channels that contribute least to the subsequent layer’s input. For the non-uniform pruning variant of our framework, we compare it with CUP. CUP assigns to each neuron/channel a feature, performs hierarchical clustering based on these features, and replaces each neuron/channel in each cluster by a cluster representative according to a maximum norm criterion. As explained, our proposed algorithm enhances all three steps of CUP. Through an ablation study, we found that each improvement is essential for achieving a competitive advantage over CUP, indicating that these enhancements complement each other. In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC, incorporating all improvements. Furthermore, the trivial baseline of random structured pruning discards neurons or channels based on a uniform probability distribution, while L1 structured pruning targets those with the smallest L1 norm of their weights or kernels.
@@
Datasets and networks. We evaluate our framework on the MNIST and CIFAR datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), LeNet (LeCun et al., 1998), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet datasets, with its performance evaluated across models such as VGG, ResNet56, and ResNet18 (He et al., 2016).
7.1 MNIST Dataset, Pairs 3-5 and 4-9
The first experiment is performed on the binary classification tasks of pairs 3/5 and 4/9 of the MNIST dataset and so we can utilize both single-output and multi-output versions of our algorithm for the compression of the final hidden linear layer. In Tables 1 and 2, we compare single-output TropNNC and TropNNC with the corresponding single-output and multi-output algorithms from Misiakos et al. (2022). We use the same network, a simple CNN network with two fully connected layers and a final hidden linear layer of size $1000$, and the same pruning ratios as in the corresponding experiment of Misiakos et al. (2022). According to Tables 1 and 2, our algorithm performs better than Misiakos et al. (2022) with particular difference in performance when using $0.3\%$ of the nodes of the hidden layer.
Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
100.0 99.42 $\pm$ 0.05 99.42 $\pm$ 0.05 99.44 $\pm$ 0.06 99.44 $\pm$ 0.06
10.0 99.38 $\pm$ 0.04 99.43 $\pm$ 0.04 99.41 $\pm$ 0.06 99.34 $\pm$ 0.16
5.0 99.37 $\pm$ 0.06 99.42 $\pm$ 0.06 99.37 $\pm$ 0.04 99.31 $\pm$ 0.16
1.0 99.33 $\pm$ 0.05 99.41 $\pm$ 0.04 99.32 $\pm$ 0.06 99.33 $\pm$ 0.05
0.5 99.12 $\pm$ 0.28 99.43 $\pm$ 0.04 99.40 $\pm$ 0.08 99.37 $\pm$ 0.04
0.3 95.98 $\pm$ 3.16 99.43 $\pm$ 0.04 98.39 $\pm$ 0.99 99.39 $\pm$ 0.10
Table 1: Comparison of methods on MNIST-3_5
Percentage of Remaining Neurons Zonotope K-means TropNNC, single output Neural Path K-means TropNNC
100.0 99.55 $\pm$ 0.04 99.55 $\pm$ 0.04 99.56 $\pm$ 0.04 99.56 $\pm$ 0.04
10.0 99.60 $\pm$ 0.07 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.05
5.0 99.56 $\pm$ 0.04 99.54 $\pm$ 0.04 99.52 $\pm$ 0.04 99.50 $\pm$ 0.03
1.0 99.35 $\pm$ 0.29 99.54 $\pm$ 0.04 99.49 $\pm$ 0.14 99.47 $\pm$ 0.11
0.5 98.03 $\pm$ 0.71 99.54 $\pm$ 0.04 99.01 $\pm$ 0.50 99.48 $\pm$ 0.12
0.3 86.38 $\pm$ 7.59 99.54 $\pm$ 0.04 94.25 $\pm$ 4.57 99.47 $\pm$ 0.09
Table 2: Comparison of methods on MNIST-4_9
@@ 7.2 MNIST and Fashion-MNIST Datasets
The second experiment is performed on the MNIST and FASHION-MNIST datasets. This classification task is multiclass, and thus only our multi-output algorithm may be used. Table 3 compares TropNNC with Neural Path K-means of Misiakos et al. (2022) for the same CNN network as above, and for the same pruning ratios as in (Misiakos et al., 2022). We again compress the final linear layer. As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to ”deepNN”, a fully connected neural network with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 7(a) and 7(b). As illustrated, TropNNC outperforms both Neural Path K-means from (Misiakos et al., 2022) and ThiNet from (Luo et al., 2017).
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to ”deepCNN2D”, a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 7(c) and 7(d). The results demonstrate that TropNNC outperforms Neural Path K-means from (Misiakos et al., 2022) and matches, or even surpasses, ThiNet from (Luo et al., 2017).
@@
7.3 CIFAR Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 8(a) and 8(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 8(c) and 8(d) show the compression of VGG’s convolutional layers for these datasets. For VGG, we applied compression to the fused networks without batch normalization, as we found that ThiNet’s performance on the original networks was practically identical to its performance on the fused networks.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means from (Misiakos et al., 2022). Furthermore, it at least matches or even surpasses the performance of ThiNet from (Luo et al., 2017) for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet. For the compression of convolutional layers of VGG, we do not present results for iterative TropNNC because it had almost identical performance to the non-iterative algorithm.
(a) AlexNet, linear, CIFAR10
(b) AlexNet, linear, CIFAR100
@@ 7.3 CIFAR Datasets
7.4 Non-ReLU Activation
In this experiment, we evaluate the performance of our algorithm in compressing networks that do not use ReLU activations, such as LeNet (LeCun et al., 1998). We conduct experiments on the MNIST and FASHION-MNIST datasets, focusing on compressing the linear and convolutional layers of LeNet. The results are presented in Figures 9(a), 9(b), 9(c) and 9(d). Interestingly, our algorithm demonstrates robust performance even on networks with non-ReLU activations.
(a) LeNet, linear, MNIST
(b) LetNet, linear, F-MNIST
(c) LeNet, conv., MNIST
(d) LetNet, conv., F-MNIST
Figure 9: Compression of LeNet on MNIST Datasets.
@@ 7.5 Non-uniform Pruning
In this section, we evaluate the effectiveness of the non-uniform variant of TropNNC. Our experiments here compress various models, such as VGG on CIFAR-10, ResNet56 on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 4, 5, 6, and 7. We should emphasize that we did not apply any form of fine-tuning or re-training. In Table 4 we present the compression of fused models, i.e. models whose batch normalization has been fused into its preceding convolutional or linear layer, as explained in a previous section. We noticed that the performance of CUP was better on the original networks with batch normalization. In Tables 5, 6 and 7 we compare our performance with the performance of CUP compressing the original networks. For our method only, we applied batch normalization fusion following the second approach described earlier—fusing the batch normalization layer with the preceding layer on a layer-by-layer basis, followed by compression of the fused layer before moving on with the next layer. We observed that our method performed significantly better when the clustering vectors were generated using the pre-fusion weights/kernels of the layers. While the reason for this improvement remains unclear, it consistently led to better performance in our experiments.
Our findings indicate that our method demonstrates a clear advantage. Notably, our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet models. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for ResNet architectures. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
@@
Model Method Threshold #params $\downarrow$ FLOPS $\downarrow$ Acc. $\uparrow$
VGG Original - 14.72M 0.63G 93.64
CIFAR10 CUP 0.15 4.70M 0.40G 56.22
TropNNC (v2) 1.12 4.70M 0.33G 90.82
ResNet56 Original - 0.85M 0.25G 93.67
CIFAR10 CUP 0.55 0.79M 0.18G 45.89
TropNNC (v2) 1.2 0.65M 0.18G 76.50
ResNet18 Original - 11.68M 3.63G 69.10
ImageNet CUP 0.6 11.62M 3.50G 58.20
TropNNC (v2) 1 11.60M 3.47G 62.60
Table 4: Comparison of CUP and TropNNC (variant 2) accuracy across different pruning thresholds for fused models without batch normalization.
@@
Original - 0.85M 0.25G 93.67
CUP 0.55 0.72M 0.20G 87.32
TropNNC (v2) 1.17 0.71M 0.19G 90.53
CUP 0.6 0.63M 0.17G 78.83
TropNNC (v2) 1.23 0.63M 0.17G 87.04
CUP 0.65 0.54M 0.15G 71.48
TropNNC (v2) 1.3 0.54M 0.15G 73.03
TropNNC (v1) 0.067 0.49M 0.19G 82.70
Table 6: Comparison of CUP and TropNNC accuracy across different pruning thresholds on CIFAR10, ResNet56.

### v3

@@ 4.2 Multi-output
7:   $(\tilde{\mathbf{a}}_{k}^{T},\tilde{b}_{k})\leftarrow\frac{\sum_{j=1}^{m}\sum_{i\in I_{k}}C_{ji}(\mathbf{a}_{i}^{T},b_{i})\widetilde{C}_{jk}}{\sum_{j=1}^{m}\widetilde{C}_{jk}^{2}}$
8:  end for
9:end for
10: Construct the new linear layer as in the non-iterative algorithm.
Proposition 6.
Suppose the clusters K are enough so that for every cluster, no two $(\mathbf{a}_{i}^{T},b_{i}),(\mathbf{a}_{i^{\prime}}^{T},b_{i^{\prime}})$ of the cluster form an obtuse angle. Then, a variant of Iterative TropNNC produces a neural network with output $\tilde{v}$ satisfying:
$\frac{1}{\rho}\max_{\mathbf{x}\in B}\|\mathbf{v}(\mathbf{x})-\tilde{\mathbf{v}}(\mathbf{x})\|_{1}\leq\\ \sqrt{m}\sum_{i=1}^{n}\min\left\{\|\mathbf{C}_{:,i}\|\|(\mathbf{a}_{i}^{T},b_{i})\|,\frac{l_{k(i)}}{N_{min}}+\|\epsilon_{:,i}\|_{F}\right\}\\ +\sum_{j=1}^{m}\sum_{i\in N_{j}}|c_{ji}|\|(\mathbf{a}_{i}^{T},b_{i})\|$
@@
Heuristic Improvements
We obtain an advantage by dropping the bias from the clustering vectors – which leads to better upper hull approximations – and normalizing the vectors used for clustering. These two heuristics are discussed in detail in the appendix.
Deep and Convolutional Networks.
For deep networks, we can view each hidden layer of the network as an instance for our algorithm, having an input from the previous layer and providing an output to the next layer. Hence, we can recursively apply our multi-output algorithm to compress each hidden layer. For convolutional networks, we unravel the kernels of the weight tensors row-wise for $\mathbf{A}$ and column-wise for $\mathbf{C}$, apply our algorithm, and reshape them back to 4D tensors. If we also have batch normalization, then this can be dealt with by fusing the operations of batch normalization into the preceding convolutional or linear layer.
Non-uniform compression.
@@
In Algorithm 2 we intentionally did not mention how clustering is performed. We can either use K-Means with a constant pruning ratio for all layers, or we can use hierarchical clustering with a global threshold parameter, like in the CUP framework. When used with hierarchical clustering, TropNNC differs from CUP in step 1, where we choose a different approach to build the clustering vectors (or filter features as Duggal et al. (2021) call them) of a convolutional layer, and more importantly in step 3, where we choose a different cluster representative based on tropical geometry. We also propose a modification to step 2 of CUP. Since the clustering vectors of different layers have varying dimensions, and vectors in higher-dimensional spaces tend to be more spread out, we introduce two variants for selecting the distance threshold for each layer:
@@
Limitations.
Limitations are discussed in the appendix.
5 Related Work
Our work falls under the category of structured pruning (He and Xiao, 2024). Here, approaches include data-free norm-based pruning (Li et al., 2017; He et al., 2019), KL-divergence based pruning (Luo and Wu, 2020), data-driven activation-based pruning (Lin et al., 2020; Sui et al., 2021), and data-driven filter pruning based on reconstruction loss minimization. Within the latter category, ThiNet (Luo et al., 2017) is a representative method that greedily removes filters contributing least to the next layer’s input, based on a data-driven reconstruction objective. Our approach resembles ThiNet in its structured layer-wise compression, but unlike ThiNet, it is data-free. Although layer-wise pruning is increasingly being replaced by global strategies – such as NISP (Yu et al., 2018) and SPvR (Hassan et al., 2025), which consider the importance of filters relative to the final layer – ThiNet remains a strong structured data-driven baseline to which we compare our method. More recently, data-free methods have been proposed. For example, CUP (Duggal et al., 2021) hierarchically clusters similar channels and selects representatives by $L^{1}$ norm. Our method builds on the same intuition of redundancy reduction via clustering, but introduces three key extensions: (i) modifying the clustering vectors, (ii) adapting the distance thresholds of hierarchical clustering to layer dimensions, and (iii) using tropical geometry to select representatives.
We built on Misiakos et al. (2022), who used a discrete form of the Hausdorff distance to geometrically bound the error between tropical polynomials, framing network reduction as a zonotope order reduction problem, and proposing a basic clustering-based compression method. Their method clusters similar neurons and replaces each cluster with the representative obtained by taking the mean of the input and output weights of the cluster. We refined their central theorem using the standard Hausdorff distance and introduce a novel algorithm that formulates compression as a simultaneous zonotope approximation problem. Our method makes a better initial estimation of the compressed network’s weights, and refines them by an iterative process. In the single-output case, our approach leads to stronger theoretical guarantees. Our method also has support for convolutional layers.
@@ 6 Experiments
Baselines.
We empirically evaluate TropNNC against baselines that perform structured pruning that do not require retraining. Specifically, as baselines we choose Neural Path K-Means by Misiakos et al. (2022), ThiNet, CUP, and the simple random and L1 structured pruning baselines. Neural Path K-Means is originally designed for linear layers only. To enable a fair comparison, we extended their approach using our proposed technique, making them applicable to convolutional layers as well. For the non-uniform variant of our framework, we compare it with CUP. As explained, our proposed algorithm enhances all three steps of CUP. In the presented experiments, we compare CUP exclusively with the fully enhanced version of TropNNC. To eliminate randomness, all random methods were performed 5 times on each model and the best performing compressed model was selected for testing. All metrics report the average over 5 repetitions of the same experiments with error bars indicating the standard deviation.
Datasets and networks.
@@
We evaluate our framework on the MNIST (Deng, 2012) and CIFAR (Krizhevsky and Hinton, 2009) datasets, testing it across various models including simple multi-layer perceptrons (MLPs), convolutional neural networks (CNNs), AlexNet (Krizhevsky et al., 2012), and VGG (Simonyan and Zisserman, 2015). The non-uniform variant of our algorithm is tested on the CIFAR and ImageNet (Deng et al., 2009) datasets, with its performance evaluated across models such as VGG, and ResNet18 (He et al., 2016). The selected datasets and models are widely used benchmarks for neural network compression, and offer diversity in their complexity. We compare the methods based on test accuracies and FLOPs reduction, following the literature.
Experimental setup.
Only a single hyper-parameter needed tuning: ThiNet’s training samples used. For each dataset, we gradually increased this number until no further improvement of ThiNet was observed. $5000$ was found to suffice for both MNIST and CIFAR. The same number can be found by using the method presented in (Luo et al., 2017) for selecting the parameter.
@@ 6.1 MNIST and Fashion-MNIST Datasets
The first experiment is performed on the MNIST and Fashion-MNIST datasets. Table 1 compares TropNNC with Neural Path K-means for the same CNN network and the same pruning ratios as in (Misiakos et al., 2022). We compress the final linear layer. As shown in the results, for both datasets, our algorithm outperforms Neural Path K-means.
To evaluate the performance of our algorithm in compressing linear layers of deeper networks, we applied TropNNC to "deepNN", a MLP with layer sizes $28\times 28,512,256,128$, and $10$. The performance plots are provided in Figures 4(a) and 4(b). As illustrated, Neural Path K-Means achieves results comparable to L1-structured pruning. By contrast, TropNNC outperforms both Neural Path K-means and ThiNet.
To assess the performance of our algorithm in compressing convolutional layers, we applied TropNNC to "deepCNN2D", a LeNet-type convolutional neural network with ReLU activations. The performance plots are provided in Figures 4(c) and 4(d). The results demonstrate that TropNNC significantly outperforms Neural Path K-means, whose effectiveness appears reduced. Moreover, TropNNC matches, or even surpasses, ThiNet.
@@
6.2 CIFAR-10 and -100 Datasets
In this experiment, we compress AlexNet and VGG trained on the CIFAR-10 and CIFAR-100 datasets to assess the performance of each compression method. Figures 5(a) and 5(b) illustrate the compression of the linear layers of AlexNet on CIFAR-10 and CIFAR-100, respectively. Additionally, Figures 5(c) and 5(d) show the compression of VGG’s convolutional layers for these datasets.
The results indicate that for larger datasets, TropNNC consistently outperforms Neural Path K-means, whose performance approaches that of random pruning. Furthermore, it matches or even surpasses the performance of ThiNet for the compression of linear layers. These findings highlight the effectiveness of TropNNC in handling more complex and larger-scale data scenarios. For the compression of convolutional layers of VGG, TropNNC matches ThiNet.
(a) AlexNet, lin., CIFAR10
(b) AlexNet, lin., CIFAR100
@@ 6.3 Non-uniform Pruning
We evaluate the effectiveness of the non-uniform variant of TropNNC. We compress various models, such as VGG on CIFAR-10, and ResNet18 on ImageNet. The results are summarized in Tables 2 and 3. We should emphasize that we did not apply any form of fine-tuning or re-training. Table 4 provides an ablation of the iterative variant of TropNNC.
Our findings indicate that our method demonstrates a clear advantage. Our approach shows a significant performance improvement for the VGG model, while the benefits are comparatively modest for ResNet18. The substantial advantage observed with VGG remains somewhat unexplained, and we have yet to determine why this effect is less pronounced for the ResNet architecture. Further investigation may be required to fully understand these discrepancies and to optimize our approach across different model types.
@@
Original - 11.69M 3.64G 69.75
CUP 0.5 11.66M 3.58G 66.80
TropNNC (v2) 1.1 11.66M 3.48G 68.95
CUP 0.6 11.49M 3.38G 53.40
TropNNC (v2) 1.2 11.46M 3.25G 61.30
CUP 0.65 11.17M 3.15G 28.45
TropNNC (v2) 1.25 10.98M 2.92G 41.65
TropNNC (v1) 0.0187 9.84M 3.46G 59.15
Table 4: Comparison of non-uniform TropNNC and iterative TropNNC (variant 1) accuracy on CIFAR10, VGG.